OpenStudy (anonymous):

solving with variables on both sides: 7m+8m = -10+5m (show checking step for the following) 5k+1 = 3k+k-3.8 12+7x = 14.2+9x+1.8

OpenStudy (anonymous):

7m + 8m = -10 + 5m 15m= -10 + 5m 10m= -10 m= -1

OpenStudy (anonymous):

5k + 1 = 3k + k -3.8 k= -4.8

OpenStudy (anonymous):

12+ 7x = 14.2 + 9x + 1.8 -4 = 2x x= -2

OpenStudy (anonymous):

@helpmehelpyou, how did you get the answer for the second question?